OpenStudy (anonymous):

find the radius of a circle whose circumference is 39.564 cm

OpenStudy (saifoo.khan):

\[C = \pi \frac{r}{2}\]

OpenStudy (anonymous):

hey saifoo can you help me after?

OpenStudy (saifoo.khan):

@HorribleAtMath , i can try for sure! ;D

OpenStudy (anonymous):

ok please go to my last question :)

OpenStudy (apoorvk):

circumference= c = 2* pi* r you have the 'c'. now substitute and 'r' post you answer and verify!